# Designing Beyond Language: Sociotechnical Barriers in AI Health Technologies for Limited English Proficiency

### v1

@@ 5.4. When AI might help, but not be the best answer
Our study showed that while AI presents promising opportunities to support LEP individuals seeking care, they may not always be the most appropriate or effective solution. In some cases, strengthening existing practices or deploying low-tech strategies may be more impactful than introducing additional technology complexity.
First, researchers generally agree that technology should aim to complement, rather than replace, human expertise in healthcare settings (Alrassi et al., 2021; Sezgin, 2023). Navigators in our study aligned with these views, consistently emphasizing the irreplaceable role of human presence (and associated trust, empathy, and relational comfort) in care.
@@
Second, communication needs could be addressed through offline resources rather than technology. Static resources (e.g., FAQs, worksheets, pamphlets) that can be helpful for conveying important health information are traditionally written by providers or navigators. Rather than attempting to replace these materials by communicating with new technology, AI could instead play a supporting role by, for example, generating these resources that are then refined by humans. These resources could then be distributed with minimal digital access as printed sheets or shared audio files (Verran, 2024), avoiding the privacy, technology access, or literacy burdens associated with AI tools. This workflow shows how AI can strengthen existing communication channels without introducing additional barriers.
@@
Third, educational scaffolding should accompany any AI interventions, as adopting new technologies often requires new skills. Navigators described investing significant effort in teaching their clients foundational skills such as using computers, sending emails, and logging into online patient portals. Introducing AI would necessitate new forms of digital literacy to ensure safe and productive interactions. In some cases, these barriers may outweigh potential benefits, suggesting that investments in improving existing workflows, interpreter access, or patient education may create more immediate impact as a first step rather than deploying new AI tools.
@@ 5.5. Limitations and Future Work
While the storyboards and scenarios provided a useful structure for discussion, they may have guided participants toward particular ideas or interpretations. Nevertheless, we found that grounding the discussion in the participants’ experiences allowed them to critically reflect on how AI might affect their existing practices and relationships. Our findings are also context-specific and do not necessarily generalize to other LEP populations or even Spanish-speaking populations in different regions of the U.S. Cultural, linguistic, and systemic characteristics of any particular area may shape the needs and experiences of those patients and navigators in ways that may differ from our study context. Future work should seek out the perspectives of LEP patients themselves and how they perceive their interactions with navigators and providers and their views on AI technology for healthcare. Building prototypes or low-fidelity design probes informed by the findings of this study, such as voice-based interfaces, could ground the participatory discussions with patients while offering insight into the validity and extensions of our findings. In this way, we can ensure that AI systems are built to be responsive and effective to the needs of the communities they intend to serve.
@@ 6. Conclusion
Appendix A Appendix
A.1. AI Literacy Questionnaire
@@
A.2. Storyboards
We include the list of all the storyboards in our study in Figure A1.
@@ A.2. Storyboards
Figure A1. Storyboards used in interviews with patient navigators.: storyboards

### v2

@@ 5.3. When AI might not be the best approach
Our study highlighted that while AI systems hold potential to support LEP individuals seeking care, they may not always be the most appropriate or effective solution. In some contexts, introducing complex technological ecosystems may create more friction than strengthening existing or deploying low-tech strategies.
Navigators in our study echoed the call in research communities for technology that complements, rather than replaces, human expertise in healthcare settings (Alrassi et al., 2021; Sezgin, 2023). However, our results highlight a tension: while most navigators were optimistic that the benefits of AI for LEP individuals would outweigh risks, successful deployment of AI places the burden on navigators to teach and monitor their clients who may be unfamiliar with or unwilling to adopt new technologies due to barriers in literacy and access to hardware. Thus, LEP adoption of AI creates a form of invisible labor (Nardi and ENgeström, 1999), which navigators expressed their willingness to shoulder because increasing digital literacy and overall education empowers patients to advocate for themselves. However, because human-in-the-loop systems would offload burdens of error checking and training on these already-stretched navigators, designing AI for this ecosystem may require intermediaries (Rahman et al., 2021) specifically for navigators.
@@
Many communication needs could also be adequately addressed through offline resources rather than technology. Static resources (e.g., FAQs, worksheets, pamphlets) that are helpful for conveying important health information are traditionally written by providers or navigators. Rather than attempting to replace these materials, AI could instead play a supporting role by, for example, generating these resources that are then refined by humans. These resources could then be distributed with minimal digital access as printed sheets or shared audio files (Verran, 2024). This hybrid workflow allows providers and navigators to leverage the generative power of AI without imposing the privacy risks and usability hurdles of LEP patient-AI interaction (Toyama, 2015).
@@
Introducing AI necessitates AI literacy that goes beyond basic digital skills (Long and Magerko, 2020). Navigators described already investing significant effort in teaching their clients foundational digital skills such as using computers, sending emails, and logging into online patient portals. Learning additional skills such as prompt engineering or information validation increases the barrier to entry for patients to adopt AI safely and productively (Xie et al., 2025). When the cognitive load of learning the tool may outweigh immediate utility, investments in improving existing workflows, interpreter access, or patient education may create more immediate impact as a first step rather than deploying new AI tools.
@@ 5.4. Limitations and Future Work
While the storyboard scenarios provided a useful structure for discussion, they reflect hypothetical scenarios, thus capturing patient navigators’ expectations and attitudes rather than observed behaviors. Furthermore, the storyboards may have guided participants toward particular ideas or interpretations. Nevertheless, because all but one of the interviewed navigators found the scenarios to be realistic, they were able to relate the fictional scenarios to their actual lived experiences, thus grounding the discussions in real examples that allowed them to critically reflect on how AI might affect their existing practices and relationships. Future work should incorporate in-situ or observational methods such as participatory design sessions or diary studies that capture actual behaviors and decision-making processes.
Our findings are also context-specific and do not necessarily generalize to other LEP populations or even Spanish-speaking populations in different regions of the U.S. Cultural, linguistic, and systemic characteristics of any particular area may shape the needs and experiences of those patients and navigators in ways that may differ from our study context. Therefore, while navigators can generalize across patients in their own regions and cultural backgrounds, they cannot speak for LEP individuals that they do not regularly interact with. Additionally, because this study captures only the perspectives of navigators, which we selected as an early-phase, risk-sensitive step into LEP communities and cultural norms before talking to patients directly, this unilaterial perspective raises two limitations. First, their position as intermediaries between patients and providers may bias their perceptions. For instance, while they may have concerns of AI around potential job displacement, all navigators emphasized aspects of their roles that were irreplaceable, such as driving patients to the hospital or teaching them how to use online portals. Nonetheless, their responses may still be inherently influenced by their positionality as navigators. Second, since these insights reflect secondhand interpretations rather than firsthand experiences of patients themselves, triangulating findings with provider and LEP patient perspectives would capture a more holistic narrative. Navigators may overestimate patients’ willingness to adopt AI tools or understate concerns that patients do not disclose to intermediaries. Therefore, future work should validate these themes through participatory sessions with LEP patients and providers. Building prototypes or low-fidelity design probes informed by the findings of this study, such as voice-based interfaces, could ground the participatory discussions with patients while offering insight into the validity and extensions of our findings. In this way, we can ensure that AI systems are built to be responsive and effective to the needs of the communities they intend to serve.
@@ 6. Conclusion
Appendix A AI Literacy Questionnaire
@@
A.1. Storyboards
We include the list of all the storyboards in our study in Figure 2.
@@ A.1. Storyboards
Figure 2. Storyboards used in interviews with patient navigators.
storyboards 6 illustrated storyboards each depicting a theme as described in Methods.